• retagged by
976 views
0 0 votes
A man alternately tosses a coin and throws a dice, beginning with the coin. Then probability that he will get a head before he gets a 5 or 6 on dice is

1) 1/4

2) 3/4

3) 4/5

4) 4/7

2 Answers

0 0 votes

since the man alternatively throws a coin and than tossess a coin where coin is tossed first..

so the sequence becomes = H,(5/6) + T,(1,2,3,4),H,(5,6) +  T,(1,2,3,4) ,T,(1,2,3,4) ,H,(5,6)..................

                                         =1/2 * 2/6 + 1/2*4/6 *1/2*2/6+................

                                        =1/6 + 1/3 *1/6 +1/3*1/3*1/6+.......

                                        =1/6( 1/(1-1/3) )= 1/4 ANSWER

0 0 votes
Correct answer is 3/4.

Probability =( prob. Of getting heads) or (prob. Of getting tail and prob. Of not getting 5,6 and then prob. Of getting heads) or (prob. Of getting tails and prob. Of not getting 5,6 and prob. Of getting tails and prob. Of not getting 5,6 and prob. Of getting heads) or ......

I think you get the idea.

So, this would be a sum of geometric progression.

Probability = 1/2 + 1/2 * 4/6 * 1/2 + 1/2 * 4/6 * 1/2 * 4/6 * 1/2 + ........

Notice how "and" & "or" in the story expression becomes * & + respectively in the mathematical expression.

Now the sum of Geometric progression with r<1 is :

Probability = 1/2 * (1/(1 - 1/2*4/6)) = 1/2 * 3/2 = 3/4.
Position:
Show:

Related questions

0 0 votes
1 answers 1 answer
532
532 views
0 0 votes
0 0 answers
308
308 views
0 0 votes
0 0 answers
479
479 views
Balaji Jegan asked Oct 23, 2018
479 views
In the text it was stated that a satellite with two uplink and one downlink slotted ALOHA channels can achieve a Downlink utilisation of 0.736, given an infinite amount o...
0 0 votes
1 1 answer
753
753 views
Balaji Jegan asked Oct 23, 2018
753 views
If mean = (3 median – mode)x, then value of x is1) 12) 23) 1/24) 3/2